OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-9, 9) and is perpendicular to -9x – 2y = 10 Please show all of your work.

OpenStudy (anonymous):

slope of the line -9x-2y=10 m = -9/2 slope of the line perpendicular to given can be calculated from the formula:  then substitute that slope in y=mx+c y = 2/9 x + c find c by substituting the point (-9,9)